# FedMeld: A Model-dispersal Federated Learning Framework for Space-ground Integrated Networks

### v2

@@ Abstract
To bridge the digital divide, the space-ground integrated networks (SGINs), which will be a key component of the six-generation (6G) mobile networks, are expected to deliver artificial intelligence (AI) services to every corner of the world. One mission of SGINs is to support federated learning (FL) at a global scale. However, existing space-ground integrated FL frameworks involve ground stations or costly inter-satellite links, entailing excessive training latency and communication costs. To overcome these limitations, we propose an infrastructure-free federated learning framework based on a model dispersal (FedMeld) strategy, which exploits periodic movement patterns and store-carry-forward capabilities of satellites to enable parameter mixing across large-scale geographical regions. We theoretically show that FedMeld leads to global model convergence and quantify the effects of round interval and mixing ratio between adjacent areas on its learning performance. Based on the theoretical results, we formulate a joint optimization problem to design the staleness control and mixing ratio (SC-MR) for minimizing the training loss. By decomposing the problem into sequential SC and MR subproblems without compromising the optimality, we derive the round interval solution in a closed form and the mixing ratio in a semi-closed form to achieve the optimal latency-accuracy tradeoff. Experiments using various datasets demonstrate that FedMeld achieves superior model accuracy while significantly reducing communication costs as compared with traditional FL schemes for SGINs.
@@ I Introduction
In this paper, to overcome all the above issues, we propose an SGI-FL framework called Federated learning with Model Dispersal (FedMeld) as our answer. Our mechanism is inspired by a key observation: Given the repetitive trajectories and store-carry-forward (SCF) capabilities of co-orbital satellites, as shown in Fig. 1, a satellite can transfer model parameters from one serving region $i$ (e.g., A) to the next region $i+1$ (e.g., B) by mixing models from these two adjacent areas; Similarly, the circular nature of satellite orbits ensures that the model parameters from region $i+1$ (e.g., B) will eventually be returned to region $i$ (e.g., A). Consequently, without requiring dedicated ground stations or ISLs, parameter mixing across all regions is naturally achieved through model dispersal via satellite trajectories, akin to how animals disperse seeds across continents. Within the FedMeld framework, users can continue training while satellites carrying model parameters traverse between adjacent regions, introducing a latency-accuracy tradeoff in managing model staleness. Moreover, model mixing ratios across regions should also be determined judiciously to achieve optimal learning performance. The core advantages of FedMeld are summarized as follows:
@@
  • FedMeld and Convergence Analysis: We propose the FedMeld framework for efficient decentralized FL in SGINs. To support effective control, we theoretically characterize the convergence bound of FedMeld under both full and partial participation scenarios, explicitly incorporating the impact of satellite constellation, inter-region model staleness, and model mixing ratio. Our analysis reveals a tradeoff between training latency and model accuracy, highlighting the need for carefully designed control strategies.
@@ III-B Convergence Analysis
Choose the learning rate $\eta_{t}=\frac{5}{\mu(\gamma+t)}$ for $\gamma=\max\left\{{\frac{{8L}}{\mu},E}\right\}-1$, when $\delta E\leq\frac{{m\left(\alpha\right){{\left({t+\gamma+1}\right)}^{2}}}}{{{{\left({t+\gamma}\right)}^{2}}+m\left(\alpha\right)\left({t+\gamma+1}\right)}}$, the convergence bound with full participation at the last time step $R$ satisfies
$\begin{split}\mathbb{E}\left[{F\left({{{\overline{\mathbf{w}}}_{R}}}\right)}\right]&-F\left({{{\mathbf{w}}^{\star}}}\right)\leq\left({\frac{1}{{1-\alpha}}-\alpha+\frac{1}{2}}\right)\frac{{L}}{{\mu\left({R+\gamma}\right)}}\\ &\cdot\left[{\frac{{25B}}{{{8\mu}}}+\frac{{\mu\left({\gamma+1}\right)}}{2}\mathbb{E}{{\left\|{{{\overline{\mathbf{w}}}_{1}}-{{\mathbf{w}}^{\star}}}\right\|}^{2}}}\right]+L{Q_{R-1}},\end{split}$ (12)
@@
$\left\{\begin{array}[]{l}{\zeta_{1}}=\frac{L}{{\mu\left({R+\gamma}\right)}}\left[{\frac{{25B}}{{{8\mu}}}+\frac{{\mu\left({\gamma+1}\right)}}{2}\mathbb{E}{{\left\|{{{\overline{\mathbf{w}}}_{1}}-{{\mathbf{w}}^{\star}}}\right\|}^{2}}}\right],\\ {\zeta_{2}}=L\left[{\frac{{1+b}}{2}\mathop{\max}\limits_{i\in\mathcal{M}}\left\{{\frac{{{N_{i}}-{U_{i}}}}{{M{U_{i}}\left({{N_{i}}-1}\right)}}}\right\}+1}\right],\\ {\zeta_{3}}=2L\left({1+\frac{1}{b}}\right){\left({E-1}\right)^{2}}{G^{2}}\eta_{R}^{2}\mathop{\max}\limits_{i\in\mathcal{M}}\left\{{\frac{{{N_{i}}-{U_{i}}}}{{M{U_{i}}\left({{N_{i}}-1}\right)}}}\right\}.\end{array}\right$.
@@ IV-C2 Optimal Solution of $\alpha$
where ${g_{1}}\left(\alpha\right)=-\rho{D_{2}}{\alpha^{2}}+\left[{{\left({{\rho^{\frac{{K+\delta}}{{K+1}}}}+\rho}\right){D_{1}}}+\left({\rho+1}\right){D_{2}}}\right]\alpha-\left({{\rho^{\frac{{K+\delta}}{{K+1}}}}{D_{1}}+{D_{2}}}\right)$ and ${g_{2}}\left(\alpha\right)={{\left({{\rho^{\frac{{K+\delta}}{{K+1}}}}+\rho}\right){\alpha^{2}}-2{\rho^{\frac{{K+\delta}}{{K+1}}}}\alpha+{\rho^{\frac{{K+\delta}}{{K+1}}}}-1}}$. Here, we use ${D_{1}}=4{\left({E-1}\right)^{2}}{G^{2}}\frac{{{\rho^{\frac{1}{{K+1}}}}}}{{{{\left({{\rho^{\frac{1}{{K+1}}}}-1}\right)}^{2}}}}$ and ${D_{2}}=4{\left({E-1}\right)^{2}}{G^{2}}\frac{{{\rho^{\frac{1}{{K+1}}}}}}{{{\rho^{\frac{1}{{K+1}}}}-1}}$ for notation simplicity. Note that $g_{1}\left(\alpha\right)$ is monotonically increasing when $\alpha$ is no more than $\frac{{\left({{\rho^{\frac{{K+\delta}}{{K+1}}}}+\rho}\right){D_{1}}+\left({\rho+1}\right){D_{2}}}}{{2\rho{D_{2}}}}$, which is less than $\frac{1}{2}$. Specifically, ${g_{1}}\left(0\right)<0$, ${g_{1}}\left({\frac{1}{2}}\right)<0$ and ${g_{1}}\left(1\right)>0$ always hold. For $g_{2}\left(\alpha\right)$, its discriminant is negative, denoting that $g_{2}\left(\alpha\right)$ is always positive. In addition, $g_{2}\left(\alpha\right)$ is monotonically decreasing when $\alpha\leq\frac{{{\rho^{\frac{{K+\delta}}{{K+1}}}}}}{{{\rho^{\frac{{K+\delta}}{{K+1}}}}+\rho}}$, where $\frac{{{\rho^{\frac{{K+\delta}}{{K+1}}}}}}{{{\rho^{\frac{{K+\delta}}{{K+1}}}}+\rho}}$ falls in the range of $\left[{\frac{1}{2},1}\right)$. The characteristics of $g_{1}\left(\alpha\right)$ and $g_{2}\left(\alpha\right)$ show that $\frac{{{g_{1}}\left(\alpha\right)}}{\left[{{g_{2}}\left(\alpha\right)}\right]^{2}}$ is monotonically increasing when $\alpha\in\left[0,\frac{1}{2}\right]$. Simultaneously, $\frac{1}{{{{\left({1-\alpha}\right)}^{2}}}}-1$} is a monotonically increasing function of $\alpha$. Thus, we can say that derivative function $f^{{}^{\prime}}\left(\alpha\right)$ is non-decreasing with respect to $\alpha\leq\frac{1}{2}$. Combining with the constraint (23), the optimal solution $\alpha^{*}$ can be determined by
@@
${\dot{\alpha}}=\left\{\begin{array}[]{l}\alpha_{0}:{f^{{}^{\prime}}}\left(\alpha_{0}\right)=0,\quad{\rm{if}}\;{f^{{}^{\prime}}}\left({\frac{1}{2}}\right)>0,\\ \frac{1}{2},\quad{\rm{if}}\;{f^{{}^{\prime}}}\left({\frac{1}{2}}\right)<0.\end{array}\right$. (26)
@@ V-A Experiment Settings
Similar to previous works [19, 20, 17, 44], we conduct image classification tasks using the CIFAR-10 [45] and MNIST [46] datasets, employing the ResNet-18 architecture [47]. Given that image classification is a key task in remote applications like remote sensing and healthcare, these datasets provide a relevant and practical basis for validating our proposed FedMeld. The CIFAR-10 dataset comprises 50,000 training samples and 10,000 testing samples across 10 categories, while MNIST contains 60,000 training samples and 10,000 testing samples with labels ranging from 0 to 9. Each client possesses an equal number of images, with distinct images allocated to each device. The necessary parameters for executing the optimization problem are estimated using the method in [48]. The computational capacity of each client is $h_{j}=15.11$ TFLOPS. In each global round, 80% of clients from each cluster are randomly selected to participate in the training process. Furthermore, we set $E=5$ [25] and $|\varsigma|=64$ [49]. The initial learning rate is determined using a grid search method [43]. The maximum tolerable delay, $T_{\max}$, is 32 hours for CIFAR-10 and 20 hours for MNIST. Additional key wireless parameters are provided in Table II [50].
@@
V-B Communication Costs

### v3

@@ Abstract
To bridge the digital divide, space-ground integrated networks (SGINs) are expected to deliver artificial intelligence (AI) services to every corner of the world. One key mission of SGINs is to support federated learning (FL) at a global scale. However, existing space-ground integrated FL frameworks involve ground stations or costly inter-satellite links, entailing excessive training latency and communication costs. To overcome these limitations, we propose an infrastructure-free federated learning framework based on a model dispersal (FedMeld) strategy, which exploits periodic movement patterns and store-carry-forward capabilities of satellites to enable parameter mixing across large-scale geographical regions. We theoretically show that FedMeld leads to global model convergence and quantify the effects of round interval and mixing ratio between adjacent areas on its learning performance. Based on the theoretical results, we formulate a joint optimization problem to design the staleness control and mixing ratio (SC-MR) for minimizing the training loss. By decomposing the problem into sequential SC and MR subproblems without compromising the optimality, we derive the round interval solution in a closed form and the mixing ratio in a semi-closed form to achieve the optimal latency-accuracy tradeoff. Experiments using various datasets demonstrate that FedMeld achieves superior model accuracy while significantly reducing communication costs as compared with traditional FL schemes for SGINs.
@@ I Introduction
To overcome all the above issues, we propose an SGI-FL framework called Federated learning with Model Dispersal (FedMeld) as our answer. Our mechanism is inspired by a key observation: Given the repetitive trajectories and store-carry-forward (SCF) capabilities of co-orbital satellites, as shown in Fig. 1, a satellite can transfer model parameters from one serving region $i$ (e.g., A) to the next region $i+1$ (e.g., B) by mixing models from these two adjacent areas; Similarly, the circular nature of satellite orbits ensures that the model parameters from region $i+1$ (e.g., B) will eventually be returned to region $i$ (e.g., A). Consequently, without requiring dedicated ground stations or ISLs, parameter mixing across all regions is naturally achieved through model dispersal via satellite trajectories, akin to how animals disperse seeds across continents. Within the FedMeld framework, users can continue training while satellites carrying model parameters traverse between adjacent regions, introducing a latency-accuracy tradeoff in managing model staleness. Moreover, model mixing ratios across regions should also be determined judiciously to achieve optimal learning performance. The core advantages of FedMeld are summarized as follows:
@@
  • FedMeld with convergence analysis: We propose the FedMeld framework for efficient decentralized FL in SGINs. To support effective control, we theoretically characterize the convergence bound of FedMeld under both full and partial participation scenarios, explicitly incorporating the impact of satellite constellation, inter-region model staleness, and model mixing ratio. Our analysis reveals a tradeoff between training latency and model accuracy, highlighting the need for carefully designed control strategies.
@@ III-B Convergence Analysis
Choose the learning rate $\eta_{t}=\frac{2}{\mu(\gamma+t)}$ for $\gamma=\max\left\{{\frac{{8L}}{\mu},E}\right\}-1$, when $\delta E\leq\frac{{m\left(\alpha\right){{\left({t+\gamma+1}\right)}^{2}}}}{{{{\left({t+\gamma}\right)}^{2}}+m\left(\alpha\right)\left({t+\gamma+1}\right)}}$, the convergence bound with full participation at the last time step $R$ satisfies
$\begin{split}\mathbb{E}\left[{F\left({{{\overline{\mathbf{w}}}_{R}}}\right)}\right]&-F\left({{{\mathbf{w}}^{\star}}}\right)\leq\left({\frac{1}{{1-\alpha}}-\alpha+\frac{1}{2}}\right)\frac{{L}}{{\mu\left({R+\gamma}\right)}}\\ &\cdot\left[{\frac{{2B}}{{{\mu}}}+\frac{{\mu\left({\gamma+1}\right)}}{2}\mathbb{E}{{\left\|{{{\overline{\mathbf{w}}}_{1}}-{{\mathbf{w}}^{\star}}}\right\|}^{2}}}\right]+L{Q_{R-1}},\end{split}$ (12)
@@
$\left\{\begin{array}[]{l}{\zeta_{1}}=\frac{L}{{\mu\left({R+\gamma}\right)}}\left[{\frac{{2B}}{{{\mu}}}+\frac{{\mu\left({\gamma+1}\right)}}{2}\mathbb{E}{{\left\|{{{\overline{\mathbf{w}}}_{1}}-{{\mathbf{w}}^{\star}}}\right\|}^{2}}}\right],\\ {\zeta_{2}}=L\left[{\frac{{1+b}}{2}\mathop{\max}\limits_{i\in\mathcal{M}}\left\{{\frac{{{N_{i}}-{U_{i}}}}{{M{U_{i}}\left({{N_{i}}-1}\right)}}}\right\}+1}\right],\\ {\zeta_{3}}=2L\left({1+\frac{1}{b}}\right){\left({E-1}\right)^{2}}{G^{2}}\eta_{R}^{2}\mathop{\max}\limits_{i\in\mathcal{M}}\left\{{\frac{{{N_{i}}-{U_{i}}}}{{M{U_{i}}\left({{N_{i}}-1}\right)}}}\right\}.\end{array}\right$.
@@ IV-C2 Optimal Solution of $\alpha$
where ${g_{1}}\left(\alpha\right)=-\rho{D_{2}}{\alpha^{2}}+\left[{{\left({{\rho^{\frac{{K+\delta^{*}}}{{K+1}}}}+\rho}\right){D_{1}}}+\left({\rho+1}\right){D_{2}}}\right]\alpha-\left({{\rho^{\frac{{K+\delta^{*}}}{{K+1}}}}{D_{1}}+{D_{2}}}\right)$ and ${g_{2}}\left(\alpha\right)={{\left({{\rho^{\frac{{K+\delta^{*}}}{{K+1}}}}+\rho}\right){\alpha^{2}}-2{\rho^{\frac{{K+\delta^{*}}}{{K+1}}}}\alpha+{\rho^{\frac{{K+\delta^{*}}}{{K+1}}}}-1}}$. Here, we use ${D_{1}}=4{\left({E-1}\right)^{2}}{G^{2}}\frac{{{\rho^{\frac{1}{{K+1}}}}}}{{{{\left({{\rho^{\frac{1}{{K+1}}}}-1}\right)}^{2}}}}$ and ${D_{2}}=4{\left({E-1}\right)^{2}}{G^{2}}\frac{{{\rho^{\frac{1}{{K+1}}}}}}{{{\rho^{\frac{1}{{K+1}}}}-1}}$ for notation simplicity. Note that $g_{1}\left(\alpha\right)$ is monotonically increasing when $\alpha$ is no more than $\frac{{\left({{\rho^{\frac{{K+\delta^{*}}}{{K+1}}}}+\rho}\right){D_{1}}+\left({\rho+1}\right){D_{2}}}}{{2\rho{D_{2}}}}$, which is less than $\frac{1}{2}$. Specifically, ${g_{1}}\left(0\right)<0$, ${g_{1}}\left({\frac{1}{2}}\right)<0$ and ${g_{1}}\left(1\right)>0$ always hold. For $g_{2}\left(\alpha\right)$, its discriminant is negative, denoting that $g_{2}\left(\alpha\right)$ is always positive. In addition, $g_{2}\left(\alpha\right)$ is monotonically decreasing when $\alpha\leq\frac{{{\rho^{\frac{{K+\delta^{*}}}{{K+1}}}}}}{{{\rho^{\frac{{K+\delta^{*}}}{{K+1}}}}+\rho}}$, where $\frac{{{\rho^{\frac{{K+\delta^{*}}}{{K+1}}}}}}{{{\rho^{\frac{{K+\delta^{*}}}{{K+1}}}}+\rho}}$ falls in the range of $\left[{\frac{1}{2},1}\right)$. The characteristics of $g_{1}\left(\alpha\right)$ and $g_{2}\left(\alpha\right)$ show that $\frac{{{g_{1}}\left(\alpha\right)}}{\left[{{g_{2}}\left(\alpha\right)}\right]^{2}}$ is monotonically increasing when $\alpha\in\left[0,\frac{1}{2}\right]$. Simultaneously, $\frac{1}{{{{\left({1-\alpha}\right)}^{2}}}}-1$} is a monotonically increasing function of $\alpha$. Thus, we can say that derivative function $f^{{}^{\prime}}\left(\alpha\right)$ is non-decreasing with respect to $\alpha\leq\frac{1}{2}$. Combining with the constraint (23), the optimal solution $\alpha^{*}$ can be determined by
@@
${\dot{\alpha}}=\left\{\begin{array}[]{l}\alpha_{0}:{f^{{}^{\prime}}}\left(\alpha_{0}\right)=0,\quad{\rm{if}}\;{f^{{}^{\prime}}}\left({\frac{1}{2}}\right)>0,\\ \frac{1}{2},\quad{\rm{if}}\;{f^{{}^{\prime}}}\left({\frac{1}{2}}\right)\leq 0.\end{array}\right$. (26)
@@ V-A Experiment Settings
Similar to previous works [19, 20, 17, 44], we conduct image classification tasks using the CIFAR-10 [45] and MNIST [46] datasets, employing the ResNet-18 architecture [47]. Given that image classification is a key task in remote applications like remote sensing and healthcare, these datasets provide a relevant and practical basis for validating our proposed FedMeld. The CIFAR-10 dataset comprises 50,000 training samples and 10,000 testing samples across 10 categories, while MNIST contains 60,000 training samples and 10,000 testing samples with labels ranging from 0 to 9. Each client possesses an equal number of images, with distinct images allocated to each device. The computational capacity of each client is $h_{j}=15.11$ TFLOPS. In each global round, 80% of clients from each cluster are randomly selected to participate in the training process. Furthermore, we set $E=5$ [25] and $|\varsigma|=64$ [48]. The initial learning rate is determined using a grid search method [43]. The maximum tolerable delay, $T_{\max}$, is 32 hours for CIFAR-10 and 20 hours for MNIST. Additional key wireless parameters are provided in Table II [49].
@@
V-B Estimation of Hyperparameters and Impact of Estimation Errors
V-B1 Estimation of Hyperparameters
First, we discuss the hyperparameters $L$ and $\mu$. Define the function $\phi_{j}\left(\mathbf{w},\mathbf{w}^{\prime}\right)=\frac{2\left[F_{j}\left(\mathbf{w}\right)-F_{j}\left(\mathbf{w}^{\prime}\right)-\left(\mathbf{w}-\mathbf{w^{\prime}}\right)^{T}\triangledown F_{j}\left(\mathbf{w}^{\prime}\right)\right]}{\left\|\mathbf{w}-\mathbf{w^{\prime}}\right\|^{2}}$. In principle, the smoothness and strong-convexity constants of each client objective $F_{j}$ are obtained as follows: $L_{j}=\sup_{\mathbf{w},\mathbf{w^{\prime}}}\phi_{j}\left(\mathbf{w},\mathbf{w}^{\prime}\right)$ and $\mu_{j}=\inf_{\mathbf{w},\mathbf{w^{\prime}}}\phi_{j}\left(\mathbf{w},\mathbf{w}^{\prime}\right)$. However, directly computing these values is infeasible for deep neural networks due to the extremely high dimensionality and the prohibitive cost of evaluating Hessian eigenvalues. Following the lightweight empirical approach in [50], we approximate these quantities using local finite-difference tests and cross-client comparisons.
First, on each device $j\in\mathcal{N}$, we perform a finite-difference test on a minibatch of size $|\varsigma_{j}|$. Given the model parameter $\mathbf{w}_{j}$, we apply a small perturbation $\varepsilon$ and compute $\phi_{j}\left(\mathbf{w}_{j}+\varepsilon,\mathbf{w}_{j}\right)$. Repeating this test for multiple perturbations $\{\varepsilon^{(m)}\}$ yields a set of empirical values $\{\phi_{j}^{(m)}\}$, where each $\phi_{j}^{(m)}=\phi_{j}\left(\mathbf{w}_{j}+\varepsilon^{(m)},\mathbf{w}_{j}\right)$ corresponds to the perturbation $\varepsilon^{(m)}$. The local smoothness estimation is obtained via a minibatch-size weighted average, which is given by $L_{\rm{local}}=\frac{\sum_{j\in\mathcal{N}}\left(\max_{m}\phi_{j}^{(m)}\right)|\varsigma_{j}|}{\sum_{j\in\mathcal{N}}|\varsigma_{j}|}$. Similarly, the local $\mu$-strongly convex constant is given by $\mu_{\rm{local}}=\frac{\sum_{j\in\mathcal{N}}\left(\min_{m}\phi_{j}^{(m)}\right)|\varsigma_{j}|}{\sum_{j\in\mathcal{N}}|\varsigma_{j}|}$.
Second, to capture the variance induced by model discrepancy across devices, we further compare client models $\left(\mathbf{w}_{j},\mathbf{w}_{p}\right)$ with $j\neq p$ evaluated on the same minibatch and compute $\phi_{j}\left(\mathbf{w}_{j},\mathbf{w}_{p}\right)$. This yields the cross-client upper and lower curvature bounds: $L_{\rm{cross}}=\max_{j\in\mathcal{N},j\neq p}\phi_{j}\left(\mathbf{w}_{j},\mathbf{w}_{p}\right)$ and $\mu_{\rm{cross}}=\min_{j\in\mathcal{N},j\neq p}\phi_{j}\left(\mathbf{w}_{j},\mathbf{w}_{p}\right)$. Finally, the global curvature estimates are taken conservatively as $\hat{L}=\max\left\{L_{\rm{local}},L_{\rm{cross}}\right\}$ and $\hat{\mu}=\min\left\{\mu_{\rm{local}},\mu_{\rm{cross}}\right\}$.
Then, we provide the estimation of stochastic gradient variance and magnitude. Similar to [43], we conduct estimation in the first global round. Each device records the stochastic gradients $\nabla F_{j}\left({\mathbf{w}}_{j};\varsigma_{j}\right)$ evaluated on its local minibatch. The empirical gradient variance and magnitude bound are given by $\hat{\sigma}_{j}^{2}={\rm{Var}}(\nabla F_{j}\left({\mathbf{w}}_{j};\varsigma_{j}\right))$ and $\hat{G}=\max_{j\in\mathcal{N}}\left\{{{{\left\|\nabla{F_{j}}\left({{\mathbf{w}}_{j};{\varsigma_{j}}}\right)\right\|}^{2}}}\right\}$.
V-B2 Impact of Estimation Errors
The empirical quantities $\hat{L}$, $\hat{\mu}$, $\hat{\sigma}_{j}^{2}$, and $\hat{G}$ obtained above are only approximations of the practical values. Here we analyze how such estimation inaccuracies influence the optimization dynamics and the resulting learning performance.
First, we discuss the relationship between the empirical estimates and the true constants. Since the empirical values $\hat{L}$ and $\hat{\mu}$ are obtained from only a few samples around certain local model points, they may underestimate the global smoothness constant and overestimate the global strong-convexity constant. As a result, it typically holds that $\hat{L}\leq L$ and $\hat{\mu}\geq\mu$. In addition, the estimation errors in $\hat{\sigma}_{j}^{2}$ and $\hat{G}$ do not exhibit a consistent directional bias, and thus do not systematically enlarge or shrink the optimization region.
Then, we discuss the influence on the optimization variables and learning performance. From (22), we observe that $\delta^{\ast}$ is independent of all hyperparameters considered in this estimation procedure. Therefore, its value is unaffected by estimation errors. However, the optimal mixing ratio $\alpha^{\ast}$ depends on the curvature-related constants through the derivative $f^{\prime}(\alpha)$ of the objective function. Let $\alpha^{\ast}$ denote the true optimal mixing ratio and $\hat{\alpha}^{\ast}$ the value obtained under estimated hyperparameters. Since the estimation biases of $\hat{\sigma}_{j}^{2}$ and $\hat{G}$ are nondirectional, we focus on the effect of the curvature constants $L$ and $\mu$. These constants influence the coefficients $\zeta_{1}$ and $\zeta_{2}$, and therefore change the shape of $f^{\prime}(\alpha)$. When $8L/{\mu}\leq E$ holds, it satisfies $8\hat{L}/{\hat{\mu}}\leq E$ and the derivative satisfies $f^{\prime}\left(\alpha\mid\hat{L},\hat{\mu}\right)<f^{\prime}\left(\alpha\mid L,\mu\right)$ for any $\alpha$. Since $f^{\prime}(\alpha)$ is non-decreasing related to $\alpha$, the root of $f^{\prime}\left(\alpha\right)$ shifts to the right, implying that $\hat{\alpha}^{*}>\alpha^{*}$ when $f^{\prime}(\frac{1}{2}\mid\hat{L},\hat{\mu})>0$. If $f^{\prime}(\frac{1}{2}\mid\hat{L},\hat{\mu})\leq 0$, the optimum remains unchanged and $\hat{\alpha}^{*}=\alpha^{*}$. On the other hand, when $8\hat{L}/{\hat{\mu}}>E$, the relationship between $\hat{\alpha}^{*}$ and $\alpha^{*}$ becomes non-monotonic and cannot be characterized analytically. When the estimated hyperparameters lead to $\hat{\alpha}^{\ast}\neq\alpha^{\ast}$, the mismatch may lead to a higher upper bound on the training loss.
Despite the estimation errors, the resulting change in the final training loss of FedMeld is modest. As demonstrated in the following experimental results, FedMeld outperforms all benchmark methods, highlighting the robustness of the proposed framework.
V-C Communication Costs